OpenStudy (anonymous):

. For any integer n,n(n2 −1)(n +2)is divisible by 4

OpenStudy (anonymous):

.....

OpenStudy (anonymous):

not a question

OpenStudy (anonymous):

for any integer n, n(n2 −1)(n +2)is divisible by 4

OpenStudy (anonymous):

does it help to know that \(n^2-1=(n+1)(n-1)\) ?

OpenStudy (anonymous):

then you have \[(n-1)\times n\times (n+1)\times (n+2)\] a product of four consecutive integers seems pretty likely that one of them has to be divisible by four

OpenStudy (anonymous):

ok yes

OpenStudy (anonymous):

thanks